A lot of mundane yet important milestones.
Dennis Wingo: We have now confirmed that the timing system on the drive for the servos and just about everything else, is back into original specifications. We found in one of the manuals that we have, the procedures for aligning the various systems of the drive. Though the other three FR-901/902’s are a little different, it is close enough that we now have verified this most critical system. We tested and confirmed that the oven controlled temperature stable clock reference (two out of three that we have tried) is within 6 x 10-8 Hertz, or better than one part in one million. This was a critical thing to find as if the system clock standards had been out of tolerance (after tweaking and 24 hour test), we would have had to have found an alternate means of providing a stable signal to the machine(s).
Fortunately as we were concerned about the possibility, we asked the ever intrepid Mark Newfield if he knew were a ultrastable 1 Mhz oscillator might be found. After a search with Ken Zin, one was located, from about the same era as the drives! We brought it over and it is hooked up and up to temperature and operating fine as a backup to the in machine standard.
Ken went through all of the rest of the timing system, and after replacing a bunch of transistors, the entire timing chain is within specification. This has had a beneficial effect on the servo system and even with the known bad bearings the servos that control the reel motors are within specification for testing at least. We still have out the reel motors and the capstan motor for bearing refurbishing. We have found more documentation in the manuals with details about the mechanical specifications of the servo system and its motors. We are anxiously awaiting the aperture cards that we are getting copied to see what other documentation that we now have from the former Ampex head of field engineering.
We have been digging into the software issues associated with the data acquisition card. At this time I can get it to digitize at 1.8 million samples per second, not good but better than before. It is going to take a lot of digging into and creating our custom software for the data acquisition and storage task. I am still looking at an alternative that I may pull the trigger on this week.
We have reorganized the work space and created an area for the computer that will be connected to the drive to reside, close enough for the hook up. With all the work that we have done and with this much verified, and knowing that the likelyhood of a tape being damaged was nil, we put a Lunar Orbiter tape on the drive Friday. There is a very good reason for this even before the system is fully back within operating specifications. The reason is that the servo system needs a tape with a 500 kHz pilot tone on it to come fully into lock. To make a new one means that the record and playback system has to be fully back into specification, which it is not as of yet.
So we took a LO tape and put it on the drive to examine the test data at the front of the LO tapes. With this we can get the best idea of how far we have to go (a ways still) before the drive is 100% back into its original specification. We will get back our rollers for the tapes this week. We can mount them and get the mechanics of the system fully back within original specification. The Capstan and Reel motors are still being worked on.
When we put the tape on, we did get video but the head that we were using is not the best one. There are still some problems in the time base corrector (for the tape, not the servo system) and other electronics but we have actually read multiburst off the oscilliscope (this does not mean that the demodulator is the correct one) from the tape. Pictures to follow and the video will be up on the facebook page in the morning.
We are also sad this week to lose our two students from San Jose state (Kenneth Willians and Austin Epps) this week as the fall term is restarting. They will be here part time (20 hours a week) to help out but their dedication and hard work this summer has really helped us with doing all the things we have had to do to make the progress that we have. Good job gentlemen! I am getting both Austin and Kenneth to talk to their professors about getting elective credit for working on the project. We did over 300 hours of this on my SEDSAT project at UAH for undergrads and would like to have the students do this here as well. I would posit at least one graduate level degree could come from this project as well, so we look forward to getting images soon!
We have also been monitoring the progress to get the $50k from Doug Comstock obligated and to get the head refurbishment going. We have found out that VMI has just had a big order come in so we might be a bit slow in getting our head back. Will know more this week
A lot of mundane yet important milestones.